# JEL: C81, O32

### v1

@@ 4.2. Stable Composition of Clinical Trials
Notes: Figure 5 illustrates the stability of the heterogeneity in the composition and quality of published clinical trials between 2010 and 2019. Panel A displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal. Panel B displays two heat maps measuring the distribution of citations received by clinical trials in each calendar year from leading journals and from all publications, respectively. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. For both panels, the sample of published clinical trials is constructed with the conservative model.
@@
These proxy variables capture slightly different conceptual objects. Interpreting each, separately, provides suggestive evidence on different aspects of the research production function. Stability in public funding suggests minimal changes, at least in the cross section, in allocation of resources to clinical trials over time. [^26] NIH RePORTER data allow us to tab total expenditures associated with grants acknowledged in these publications. These grant-based expenditure measures, too, are stable. Note, however, that linking research grant dollars to specific scientific papers is an imprecise exercise. See Li, (2017) for an extended discussion of this disconnect. Stability in citation patterns, in contrast, suggests that the “usefulness” of information produced by trials also remains stable.
@@
Panel B of Figure 5 provides a more granular view of this heterogeneity. For each publication in our census, we collect citations from publications in leading journals and from all of PubMed. We plot the total (three-year) citation counts to publications at each citation quantile over time. [^27] Appendix C presents these results using alternative cuts of our data and five-year citations. The patterns remain unchanged. Grey areas at the bottom of each plot represent publications that receive zero citations. This distribution is stable over our time period. [^28] Reassuringly, Panel B indicates that this exercise does capture large shifts in the research environment: citation counts appear especially high in the later years of data, when three-year citation counts (beginning for year 2017) capture research written during the COVID-19 pandemic. Specifically, we document that the tails of the citation distribution are unchanging. Observe that we elongate the y-axis, to highlight the right-tail. Citations received at the top of this distribution are constant across periods. The left-tail is similarly stable: the share of records receiving zero citations is essentially constant, year-on-year.
@@
This result is consistent with the idea that, for clinical trials, the right-hand side of the production function, (2.1), in our simple framework—i.e., $\alpha_{t}S_{t}$—is both heterogeneous and stable over time. For this pattern to be consistent with a systematic decrease in the research productivity of clinical trials, it must be the case that there is an exactly offsetting increase in the distribution of research capital. Although we cannot reject that there are shifts in upstream, basic inputs into clinical trials or large changes in factor prices during this period, there is little credible evidence consistent with either during this time period. A small number of papers examine frictions in markets for basic research. See, for example, Hill and Stein, (2021) and Myers, (2020). To our knowledge, this literature has not provided evidence of substantial shifts in these frictions over time, which in turn would increase the cost or decrease the productivity of basic science. [^29] Anecdotal evidence suggests that there may have been substantial decreases in the cost of basic science during this time period. See Scannell et al., (2012). On factor prices, as discussed in Section 3, nearly all estimates of research and development costs are self-reported by the pharmaceutical industry. Any measured increases in these expenditures that cannot be independently verified should be viewed with caution, especially in light of our findings of stability.
More generally, any type of cancellation—an argument against the conclusion that research productivity across clinical trials is stable—would appear to be a knife-edge case. In our view, it is far more likely that the distributions of both research productivity and research capital within clinical trials have been constant over this time period.
5. Discussion
All research investments are not alike. Certain investments are more likely to yield innovations that have welfare-relevant impacts. Others, while potentially valuable contributions to a research ecosystem, may have insignificant impacts on growth. We argue that this type of heterogeneity is quantitatively significant in medical research. We use an open-source, fine-tuned large language model to construct a novel census of clinical trials from 2010 to the present. We contrast stable trends in clinical trials—in quantity, quality, and composition—with substantial growth in the quantity of other forms of medical research. Our findings rationalize, in this field, two stylized facts about research investments and innovative output: the quantity of research inputs has grown substantially, while improvements in mortality continue at the same stable rate.
@@ 5. Discussion
A second dimension of heterogeneity is helpful in contextualizing our findings: geography. Figure 6 displays trends in the annual quantity of published clinical trials, and papers citing clinical trials, disaggregated by the location of the first-listed author. [^30] In most domains of scientific and medical publication, last-listed authors are senior investigators. First-listed authors are typically junior investigators. Agha and Molitor, (2018) observe that, for large-scale clinical trials, first-authors are more likely to be the principal investigator. In our context, the first- and last-author have the same listed country for 85 percent of clinical trial records and 91 percent of non-clinical trial records. We collect details on author location from Clarivate Analytics’ Web of Science. See Section A.4 for further details. In our sample, the top four producers of published clinical trials, and papers citing clinical trials, are, in order, the United States, China, Germany, and Japan. Clinical trial investment—across geographies—is stable. [^31] Note that Panel A of Figure 6 displays measurements of the location of clinical trial publication authors, not the location of clinical trial sites. We cannot rule out that, for example, trials authored by researchers with United States mailing addresses were conducted elsewhere. Thus, we do not interpret these as facts on clinical trial “offshoring.” See Petryna, (2007) and Durvasula, (2023) for longer discussions of the geography of clinical trial sites. Roughly 30 percent of published clinical trials originate in the United States. This share is unchanging across periods. By contrast, Panel B indicates a large shift in the medical research ecosystem. Since 2013, there has been a small increase in the quantity of other medical research published by authors in the United States. This small change is dwarfed by increases—on the order of 30-50 percent—in the quantity of other medical research from China and from countries outside of the top four producers of medical research.
Figure 6. Composition of Medical Research Across Countries
@@
Notes: Figure 6 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, by country. The sample of published clinical trials is constructed with the conservative model. To address truncation, we report the number of publications that cite clinical trials published in the preceding three years.
These compositional facts are, in many ways, unsurprising. In recent years, editorials in prominent journals in medicine have observed that changes in the incentive structures for physicians, scientists, and even medical students, across countries, have led to “an extreme of quantity at the expense of quality” in medical publishing (Siegel et al.,, 2018). [^32] See, as one example, program certification requirements, which describe publication expectations for medical students and clinical faculty, from the Accreditation Council for Graduate Medical Education in Accreditation Council for Graduate Medical Education, (2017). Recent work documents a substantial increase in the quantity of meta-analyses and systematic reviews—summaries of existing clinical trials—driven by scientists from outside of the United States (Ioannidis, 2016a, ). In parallel, there has been a rapid increase in the number, and scale, of “mega-journals”—scientific journals published frequently, with a large and growing number of pages (Ioannidis et al.,, 2023).
@@
Nonetheless, substantial constraints—financial, ethical, and practical—keep firms, physicians, and scientists from increasing the frequency of clinical trials. Clinical trials are infrastructure-intensive. Trials are run at specific sites, which must identify patients who satisfy all relevant eligibility criteria and monitor them for periods ranging from six months to twenty years (Piantadosi,, 2024). Best available estimates suggest that trials can cost on the order of 500 million to one billion dollars to complete (Qiao et al.,, 2019). Perhaps most importantly, sites must be capable of producing data sufficient to persuade regulators. Certain regulators require that sponsors of new drug applications submit trial evidence collected from a domestic population. [^33] Until December 2023, Japan required domestic phase I clinical trials of drugs developed overseas before Japanese individuals could participate in international phase III trials for pharmaceutical regulatory approval. This policy reflected, in part, differences in disease burden in Japan, relative to countries where drugs are routinely tested. This may explain why Japan is a top producer of clinical trials over our sample period. See Namba et al., (2024) for one discussion. In recent years, however, the U.S. Food and Drug Administration (FDA)—responding to concerns from patient groups—has indicated a limited willingness to approve drugs on the basis of evidence collected from exclusively foreign sites (for a longer discussion, see Alsan et al.,, 2024). In one especially high-profile example, a cancer drug was rejected by the FDA after being tested exclusively in China (Kolata,, 2022). As long as the U.S. market continues to have outsized value for pharmaceutical firms—a consequence of the especially high prices paid by American consumers—it may be unsurprising to find that evidence disproportionately originates in the United States. Observational studies, meta-analyses, and case reports are, by contrast, cheap and not subject to the same financial and regulatory pressures. Digital tools, including search engines and language models, will likely further depress these costs.

### v2

@@ 4.2. Clinical Trials as a Productivity Indicator
Notes: Figure 4.2 displays three time series: the proportion of clinical trials that are publicly funded, the proportion of clinical trials that are ever cited by a leading journal, and the proportion of publicly funded clinical trials that are ever cited by a leading journal. The sample of published clinical trials is constructed with the conservative model.
We designate a publication as having public funding if it appears in the National Institutes of Health RePORTER data, as linked to a funded research grant, or if PubMed reports a source of U.S.-based research funding. [^42] Technically, PubMed indicates whether a publication received research funding from government agencies and private philanthropies outside of the United States. We restrict consideration to those sources of funding based in the United States because there appears to be a shift in the reporting of certain sources of non-U.S. funding beginning in 2017. The total number of publications linked to a source of public funding increases sharply—by roughly 50 percent from 2016—in 2017. On inspection, other trends are smooth during this time, including trends in public funding for the United States. We are unable to determine if certain sets of authors began to disclose sources of funding in 2017, if certain journals changed their reporting requirements, or if PubMed changed its data construction process. We thus focus on the United States, where we find no evidence of large changes in data reporting practices. To designate a set of the “top” journals in medicine, we follow Angrist et al., (2020). We collect all citations originating from the “trunk journals” in medicine, the Journal of the American Medical Association and the New England Journal of Medicine, to records in PubMed between 2010 and 2022. We designate a journal as being “leading” if it received at least 100 citations from a trunk journal over this time period. This yields a list of 84 journals. [^43] See Section A.3 for details. Appendix D reproduces this figure using five-year citations. The trend is unchanged.
Two facts are worth highlighting. First, all three measures are stable, essentially unchanging, over our time period. Second, nearly 70 percent of clinical trials are never cited by a leading journal, suggesting substantial, but time-invariant, heterogeneity across trials.
@@
These proxy variables capture slightly different conceptual objects, which are helpful in understanding different aspects of the production function (2.1). Stability in public funding suggests minimal changes, at least in the cross section, in allocation of resources to clinical trials over time. [^44] NIH RePORTER data allow us to tab total expenditures associated with grants acknowledged in these publications. These grant-based expenditure measures, too, are stable. Note, however, that linking research grant dollars to specific scientific papers is an imprecise exercise. See Li, (2017) for an extended discussion of this measurement challenge. Stability in citation patterns, in contrast, suggests that the “usefulness” of information produced by trials has been stable across this period.
@@
Figure 4.3 provides a more granular view of this heterogeneity. We plot quantiles of the distribution of the number of (three-year) citations to publications in our census from publications in leading journals, for each year between 2010 and 2019.454545Appendix D presents these results using alternative cuts of our data and five-year citations. The patterns remain unchanged. Grey areas at the bottom of each plot represent publications that receive zero citations. This distribution is stable over our time period. Observe that we elongate the y-axis to highlight the right-tail of the citation distribution. The quantiles corresponding to the right tail of the citation distribution are constant across years. The left-tail is similarly stable: the share of records receiving zero citations is essentially constant, year-on-year.
Figure 4.3. Citation Distribution Across Time, Clinical Trials
Notes: Figure 4.3 displays a heat map measuring the quantiles of the distribution of citations received by clinical trials in each calendar year from leading journals. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. The sample of published clinical trials is constructed with the conservative model.
Figure D.4 is constructed analogously, where instead we display quantiles of the distribution of the number of citations received from all publications indexed in PubMed. There, we find that the right-tail of this distribution is increasing over time. In other words, as the total number of publications is increasing, the total number of citations received by clinical trials is increasing. However, these citations are not originating in leading journals.
@@
We interpret this set of facts as evidence that our measured quantity of clinical trials captures an object that, though heterogeneous, is stable over time. Thus, suggestively, research productivity is similarly stable. We cannot reject that there have been shifts in upstream, basic inputs into clinical trials or changes in factor prices during this period, which might suggest that the underlying quantity of resources directed toward research is, in fact, changing over time. In our view, however, there is little rigorous evidence that supports such an argument. A small number of papers examine frictions in markets for basic research. See, for example, Hill and Stein, (2021) and Myers, (2020). To our knowledge, this literature has not provided evidence of substantial shifts in inputs or factor prices over time, which in turn would increase the cost or decrease the productivity of basic science. [^46] Anecdotal evidence suggests that there may have been substantial decreases in the cost of basic science during this time period, which would suggest the opposite conclusion in our simple framework—that research might be more productive over this period. See Scannell et al., (2012). On factor prices, as discussed in Section 3, nearly all estimates of research and development costs are self-reported by the pharmaceutical industry. Any measured increases in these expenditures that cannot be independently verified should be viewed with caution, especially in light of our findings of stability. [^47] Herper, (2022) collects anecdotal evidence, which has been widely cited, on sharply rising costs of clinical trials. Although the magnitudes included in this discussion are striking, they capture the impressions of a small number of industry stakeholders and, to our knowledge, have not been verified in a more detailed analysis.
For clinical trials to be an informative proxy for the quantity of medical research—and for the black line in Figure 4.1, then, to capture the relevant measure of inputs—we must make two additional assumptions. First, the existence of publication bias means that scientific publications do not perfectly capture research inputs (Andrews and Kasy,, 2019). In our setting—where the object of interest is the trend in research over time—we need only make a relatively weak assumption about the extent of publication bias: we assume that changes in publication bias, year-on-year, do not perfectly offset changes in the quantity of actual research.
Second, we must assume that the number of publications associated with each clinical trial is constant across our period of interest (i.e., that the amount of research “embedded” in each publication is not systematically changing over time). We view this assumption as reasonable for three reasons. First, by 2010, nearly all scientific journals in our sample had adopted rules requiring authors of clinical trial publications to adhere to a standardized format when reporting trial results (Schulz et al.,, 2010). Under these rules, the structure of a paper and its contents are pre-determined; there is little scope for authors to, for example, split results across several publications. Second, when constructing our sample of clinical trials, we have aimed to exclude those publications that re-analyze data generated by an existing study. Third, more suggestively, we use data on clinical trial publications supported by National Institutes of Health grants to examine the quantity of grant dollars per reported clinical trial publication over this period. We find no evidence of a change in this relationship between funding and publication during our period of interest. [^48] We inspect this relationship by collecting records from the NIH RePORTER database, for years 2010 to 2019. Specifically, we collect records of the total award associated with funded projects and the quantity of any publications flagged by our procedure as clinical trials. For the median NIH grant, we find that an award of roughly \$100,000 per clinical trial, and a mean of roughly \$200,000 per trial. This is stable across years of data.
4.3. Why is the Quantity of Research Rising?
The quantity of papers citing clinical trials has increased, per Figure 4.1, by nearly a factor of two over our period of interest. In this section, we examine various measures of publication composition and quality to shed light on the source of this increase. Our aim, in doing so, is to determine whether this measured increase corresponds to a “real” increase in the quantity of underlying research or whether it captures changes in the ease and frequency of publication. Evidence consistent with the latter suggests, then, that counts of scientific publications may provide little insight on questions concerning research productivity.
4.3.1. Publication Content
We inspect the technical content of “non-clinical-trial” papers in three ways. We begin by hand-labeling random samples of 100 records, drawn from the beginning (publication year 2013) and end (publication year 2019) of our sample of papers that cite clinical trials. In this pooled set of 200 records, we identify one record that should have been in our sample of clinical trials. [^49] The erroneously labelled publication is PMID 23797691, which reports the results of a clinical trial studying a drug to treat HIV. Note that this exercise suggests a false negative rate of 0.005, consistent with estimates reported previously. The remainder of the records can be split into seven sets: editorial records (e.g., comments on other papers, errata, editorials, 2.5%), publications reporting preclinical studies (e.g., in-laboratory and in-animal tests, 21.5%), clinical trial protocols (1%), randomized trials studying something other than a drug (2.5%), case studies and case reports (4.5%), observational studies (30.5%), and review articles (e.g., meta-analyses and literature reviews, 37%). [^50] In three cases, review articles appeared with the heading “Comment” or “Editorial.” We categorize these records as falling into the “Review” category. We find a 30 percent increase in the total number of review articles between 2013 and 2019, while the quantity of observational studies remains essentially constant (31 in 2013, and 30 in 2019). The increase in the quantity of meta-analyses is consistent with increases documented elsewhere (Ioannidis et al.,, 2013). Of course, in our small sample of records, it is difficult to draw authoritative conclusions about the role of growth in this type of literature.
We collect measures of changes in the quantity of meta-analyses, based on the hypothesis that this may be a quantitatively important driver of the changes in our sample, in two ways. First, we implement a simple search of the abstracts of publications in this “citing” sample. We inspect the set of hand-labeled records, used as the “training data” for our fine-tuning procedure in Section 3.3, as many were flagged as out-of-sample because they were literature reviews or meta-analyses. From this set, we collect a list of keywords that were indicators when we hand-labeled these records that the publication was a meta-analysis. Our search terms are intentionally specific, as the aim of this procedure is to yield a plausible lower bound for any trend. [^51] We searched for the following keywords: meta-analysis, metaanalysis, metaanalyses, systematic review, systematic reviews, systematically review, systematic search, review of published data, literature review, literature search, search of databases, review all literature, reviewed all literature, narrative review, systemic review. We also included any record that mentioned at least two of the following databases, which—based on our hand-labeling experience—are often referenced as part of the methods involved in a literature review or meta-analysis: MEDLINE, EMBASE, CINAHL, PubMed, Cochrane Central Register of Controlled Trials, BioMedCentral. More liberal search terms (e.g., “review”) risk capturing a large number of extraneous records. As expected, this yields a small number of records: in our three-year citing trials sample, there are 1,606 record in our citing sample that satisfy these criteria in 2013 and 3,372 in 2019, corresponding to a 109 percent increase.
Our second method generates—as the discussion in Section 3 suggests—what might be viewed as an upper bound on the quantity of meta-analyses and literature reviews in our data. We collect all records indexed in PubMed with the NLM tags “meta-analysis,” “systematic review” and “review.” There are 9,407 records in our citing sample with one of these tags in 2013 and 14,991 in 2019, corresponding to a 59 percent increase. [^52] Figure D.8 displays a time series of the number of ‘citing’ papers that satisfy both our keyword and NLM-tag based methods for identifying meta-analyses.
From these three procedures, we draw a suggestive conclusion: there has been a quantitatively important increase in the number of meta-analyses over our period of interest. This increase is on the order of 30 to 100 percent, but various measurement strategies yield alternative estimates. The specific quantity of meta-analyses is not of central interest for this paper, except insofar as it allows us to conclude that the quantity of medical research that summarizes existing papers—rather than producing new evidence—appears to be growing over time. This trend provides support for our decision to focus on clinical trials as a measure of productivity-relevant research and sheds some light on the source of documented increases in the quantity of research in other papers. As we note in Section 2, meta-analyses and similar forms of literature review can provide important information to physicians and can be vital in the diffusion of existing information, especially as the stock of knowledge grows (Jones,, 2009). Here, then, the distinction between those forms of scientific effort that are relevant to productivity metrics and those that are relevant to more expansive welfare metrics is especially important.
4.3.2. Publication Geography
Changes in the production of meta-analyses have been linked (e.g., Ioannidis et al.,, 2013) to shifts in the geographic distribution of medical research. Publication data allow us to inspect geographic trends in our sample.
@@ 4.3.2. Publication Geography
Figure 4.4 displays trends in the annual quantity of published clinical trials, and papers citing clinical trials, disaggregated by the location of the first-listed author. [^53] In most domains of scientific and medical publication, last-listed authors are senior investigators. First-listed authors are typically junior investigators. Agha and Molitor, (2018) observe that, for large-scale clinical trials, first-authors are more likely to be the principal investigator. In our context, the first- and last-author have the same listed country for 85 percent of clinical trial records and 91 percent of non-clinical trial records. We collect details on author location from Clarivate Analytics’ Web of Science. See Section A.4 for further details. In our sample, the top four producers of published clinical trials, and papers citing clinical trials, are, in order, the United States, China, Germany, and Japan.
Figure 4.4. Composition of Medical Research Across Countries
@@
Notes: Figure 4.4 displays measurements of the number of clinical trials, and papers that cite clinical trials, published in each calendar year, by country. The sample of published clinical trials is constructed with the conservative model. To address truncation, we report the number of publications that cite clinical trials published in the preceding three years.
Clinical trial quantities—across geographies—are stable over our period of interest. [^54] Note that Panel A of Figure 4.4 displays measurements of the location of clinical trial publication authors, not the location of clinical trial sites. We cannot rule out that, for example, trials authored by researchers with United States mailing addresses were conducted elsewhere. Thus, we do not interpret these as facts on clinical trial “offshoring.” See Petryna, (2007) and Durvasula, (2023) for longer discussions of the geography of clinical trial sites. Roughly 30 percent of published clinical trials originate in the United States. This share is unchanging across periods. By contrast, Panel B indicates a large shift in the medical research ecosystem. Since 2013, there has been a small increase in the quantity of other medical research published by authors in the United States. This small change is dwarfed by increases—on the order of 30-50 percent—in the quantity of other medical research from China and from countries outside of the top four producers of medical research.
@@
Institutional details are helpful in making sense of the differential patterns across these cuts of data. Regulatory requirements and local capabilities are key factors influencing where clinical trials are conducted and published. Sites must be capable of producing data sufficient to persuade regulators in high-value markets. Certain regulators require that sponsors of new drug applications submit trial evidence collected from a domestic population. [^55] Until December 2023, Japan required domestic phase I clinical trials of drugs developed overseas before Japanese individuals could participate in international phase III trials for pharmaceutical regulatory approval. This policy reflected, in part, differences in disease burden in Japan, relative to countries where drugs are routinely tested. This may explain why Japan is a top producer of clinical trials over our sample period. See Namba et al., (2024) for one discussion. In recent years, the U.S. Food and Drug Administration (FDA)—responding to concerns from patient groups—has indicated that it may not be willing to approve drugs on the basis of evidence collected from exclusively foreign sites (for a longer discussion, see Alsan et al.,, 2024). In one especially high-profile example, a cancer drug was rejected by the FDA after being tested exclusively in China (Kolata,, 2022). As long as the U.S. market continues to have outsized value for pharmaceutical firms—a consequence of the especially high prices paid by American consumers—it may be unsurprising to find that trial evidence disproportionately originates in the United States, even as global production of other forms of medical research continues to grow. Observational studies, meta-analyses, and case reports are, by contrast, cheap and not subject to the same financial and regulatory pressures. Digital tools, including search engines and the large language models deployed in this paper, will likely further depress these costs, potentially further widening this gap.
Four countries are the most prolific producers of clinical trials and other forms of medical research. Of interest here, however, are those countries that experienced the largest growth in their production of medical research over this period. On this measure, China, Poland, Japan, and Spain top the ranking in our data. In Section D.2, we provide more details on the set of countries with the largest growth in this form of medical research over our period of interest. Several details are worth noting. First, Chinese production of records in our citing sample increases by roughly 225 percent, producing 5,620 papers in our citing set in 2019. The next country in this ranking, Poland, experiences 96 percent growth and produces 311 papers in our citing set in 2019. The top 23 countries in terms of growth, listed in the appendix, produce 93 percent of all 2019 papers in this set. Per the World Bank income classification scheme, nearly every country in this fast-growing group is high income. The only countries categorized differently are China, India, Brazil, Iran, and Turkey, which the World Bank classifies as either upper (China, Brazil, Iran, Turkey) or lower (India) middle-income. [^56] https://datatopics.worldbank.org/world-development-indicators/the-world-by-income-and-region.html That is, this trend captures, in part, an increase in the quantity of diffusion-focused literature in high-income countries that have long had robust health infrastructure. Below, we establish that on several quality measures, these publications appear to be declining in value over time—consistent with the idea that some of this increase in production may be a response to changing scientific incentives for publication, rather than “real” invention or diffusion effort.
4.3.3. Publication Quality
Next, we collect a set of facts about the quality of this growing set of “citing” papers. We begin in Figure 4.5 by replicating Figure 4.3 for the sample of papers that cite clinical trials. There are several elements of this plot that are informative. First, the share of “citing” papers with a small number of citations appears to grow over time. That is—the left tail of the citation distribution for this set is both long and growing. Note also that the right tail of the citation distribution appears to shrink over time. That is, the proportion of highly cited papers appears to be shrinking. These patterns provide one indication that the central tendency of the quality of non-clinical-trial medical research may be decreasing over time. [^57] Figure D.5 gives an analogous figure, displaying the distribution in citations from all publications to the sample of publications that cite clinical trials. Here, the right tail of the distribution of citations is growing.
Figure 4.5. Citation Distribution Across Time, Papers that Cite Clinical Trials
Notes: Figure 4.5 displays a heat map measuring the quantiles of the distribution of citations received by papers that cite clinical trials, in each calendar year, from leading journals. The $y$-axis has been stretched to elongate the right-tail of the citation distribution. Colors are displayed in a log scale. The sample of published clinical trials is constructed with the conservative model. The sample of papers that cite clinical trials is constructed using a three-year window.
Citations are imperfect measures of the ex post quality of a scientific publication. Our data allow us to examine a variety of other measures that provide insight into the ex ante quality of a particular instance of research. Unfortunately, shifting geographic trends in this sample and the low quality of measures of public funding for countries outside of the United States render it difficult to use a measure of public funding as a proxy for quality. We do find a 10 percentage point decrease in the share of publications in this set that record having received (U.S.-based) public funding, but this is likely best interpreted as a mechanical consequence of the geographic heterogeneity displayed in Figure 4.4.
Instead, we focus on author characteristics, constructed using publication data, that also shed some light on research quality. Scientific author norms typically mean that first-listed authors are junior scientists and last-listed authors are senior, lead scientists. We collect facts about these groups separately.
We begin by examining the productivity of senior scientists who authored papers in our citing sample. We collect all publications indexed in PubMed authored by lead scientists (last-authors) who write at least one paper in our citing sample. The median scientist in this set writes five publications per year. We observe opposing trends in annual publication quantity and citation patterns during this time. We find that the median quantity of papers written by a scientist per year increases from 3 to 5 over this period. The median number of citations to those papers from leading journals drops, over the same period, from 3 to 1. At the 90th percentile, the quantity of per scientist publications per year is stable, at 9 per scientist between 2013 and 2019. Citations from leading journals are similarly stable, at 10 per publication. [^58] We observe sharp declines in the likelihood that a scientist has ever received funding from the NIH or from another U.S. federal agency. As noted above, however, we hesitate to place much weight on this fact, as it likely reflects changes in the composition of the sample. For the median senior scientist who appears in our citing sample in 2013, approximately 80 percent of their previously published papers (i.e., published before 2013) had received at least one citation. For scientists who first appear in our citing sample in 2019, the share of previous publications with a non-zero number of citations falls to roughly 70 percent.
For junior scientists, we focus on a measure of attrition: after authoring a paper in our citing sample, do we ever observe another publication written by that junior scientist? We collect all publication records in the Web of Science associated with these junior (first-listed) authors. In 2013, the likelihood that a junior scientist published another paper in 2014 was roughly 15 percent. By 2019, this falls to roughly 8 percent. We observe similar trends when two- and three-year publication patterns are observed. We interpret this attrition as evidence that junior scientists who author papers in our citing sample may be less likely to remain research productive in the future.
Taken together, we interpret these facts about scientist research productivity as suggestive evidence of declining average research quality in our citing sample over time. When interpreted in light of the citation distribution trends in Figure 4.5, we conclude that measured increases in the quantity of this “other” medical research may be accompanied by a decline in its average quality. It is precisely this shift that suggests that quantities of scientific publications in this sector, when not restricted to more precise categories of research such as clinical trials, may be a poor indicator of research productivity.
5. Discussion
Studies of innovation and productivity have long grappled with the prospect of diminishing returns to research (e.g., Merton,, 1935; Stafford,, 1952; Schmookler,, 1952, 1954, 1966; Griliches,, 1990, 1994; Bloom et al.,, 2020). Answers to the question at the center of this literature—is the cost of sustained technological progress increasing over time?—have been unsatisfying, as fundamental challenges in the measurement of technological innovation have limited our ability to do more than determine that available metrics do not yield conclusive answers. In one sense, this paper fits in this tradition, in that we find that, in the context of medical research, recent concerns about declining productivity are an artifact of long-standing measurement challenges.
On the other hand, enabled by advances in generative AI, we introduce a new technique for annotating unstructured text that allows us to construct a proxy for medical research that, we argue, closely approximates the object of interest in an analysis of medical research productivity. We fine-tune an open-source large language model, which allows us to extract records of clinical trials from a corpus of scientific publications. The resulting census corrects for errors in previous measures of clinical trial research and approximates the quality of hand-labeled data. We show that trends in the quantity, quality, and composition of clinical trials have been stable since 2010.
Our data allow us to contrast these trends with analogous patterns in the production of other forms of medical research. We find that, since 2010, the quantity of other forms of medical research has roughly doubled, a pattern consistent with estimates reported in other papers. On inspection, the increase in quantity includes a large increase in the number of papers authored by scientists in China and in the number of papers that synthesize existing literature (e.g., literature reviews and meta-analyses). Growth in these forms of research coincides with an average decline in various measures of publication quality. Taken together, we interpret these facts as providing no support for claims of declining research productivity, at least since 2010.
These patterns are, in many ways, unsurprising. In recent years, editorials in prominent medical journals (Harvey,, 2020) have observed that changes in the incentive structures for physicians, scientists, and even medical students, across countries, have led to “an extreme of quantity at the expense of quality” in medical publishing (Siegel et al.,, 2018). [^59] See, as one example of how these incentives have shifted, program certification requirements from the Accreditation Council for Graduate Medical Education in Accreditation Council for Graduate Medical Education, (2017). The ACGME provides formal accreditation for training programs and, as part of this process, lays out expectations of minimum quantities of publication for medical students and clinical faculty. Note that certain researchers have put forth alternative explanations for the rapid increase in the total number of publications. A trend toward evidence-based decision-making in medicine has increased demand for articles that synthesize large bodies of research (Lohr,, 2004). Recent work documents a substantial increase in the quantity of meta-analyses and systematic reviews—summaries of existing clinical trials—driven by scientists outside of the United States (Ioannidis, 2016a, ). In parallel, there has been a rapid increase in the number, and scale, of “mega-journals”—scientific journals published frequently, with a large and growing number of pages (Ioannidis et al.,, 2023).
Despite these changing incentives, substantial constraints—financial, ethical, and practical—keep firms, physicians, and scientists from increasing the quantity of clinical trials. Clinical trials are infrastructure-intensive. Trials are run at specific sites, which must identify patients who satisfy all relevant eligibility criteria and monitor them for periods ranging from six months to twenty years (Piantadosi,, 2024). Sites may struggle to find patients who meet these criteria, especially in small disease markets (Kolata,, 2017), and the costs of accruing large samples for certain studies can be prohibitive (Alsan et al.,, 2024).